OpenStudy (anonymous):

√3:(y^10) √3:(81y^11) Multiply and simplify

OpenStudy (agreene):

\[\large\sqrt[3]{y^{10}}\sqrt[3]{81y^{11}}\] Is this the question?

OpenStudy (anonymous):

Yes I am sorry but I do not know how to get those marks on my computer and I am having a really difficult time understanding this even after I am reading my text book :)

OpenStudy (anonymous):

\[3\sqrt[3]{3}y^{7}\]

OpenStudy (agreene):

Well, we can re-write this as: \[(y^{10})^{1/3}(81y^{11})^{1/3}=(y^{10})^{1/3}(81)^{1/3}(y^{11})^{1/3}=(y^{10})^{1/3}3(3^{1/3})(y^{11})^{1/3}\] Or: \[\large3\sqrt[3]3 \sqrt[3]{y^{10}}\sqrt[3]{y^{11}}\] Now, more than likely your book is going to lie and say this can be reduced further... it actually cannot, if x is negative... but they don't like teaching people that... \[\large 3\sqrt[3]3 * y^7\]

OpenStudy (anonymous):

thank u all very much